Would you pls help me to solve this:
Explain whether or not you agree or disagree with the following statement:
"Whenever you change the slope of a line, you change the intercepts of that line. It follows, then, that whenever you change intercepts of a line, you change the slope of that line."
Thanks.

Answer by fcabanski(1391)

The slope intercept form of a line, y=mx+b, provides the slope (that's m) and the y-intercept (b, the y value when x = 0).


The question suggests that changing the slope of a line changes the intercepts.


Let's try to disprove that. y=x is a line with slope 1 and y-intercept 0 (y=1z+0). If the slope changes does the intercept change? Not necessarily. y=3x is a line with slope 3, but the y intercept is still 0 (y=3x+0). Changing the slope didn't change the intercepts.


Does changing the intercepts change the slope? Not necessarily. y=3x+2 has a slope=3 and y-intercept =2. Change the intercept with y=3x+3. The slope=3, but the intercept=3. Changing the intercept didn't change the slope.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)